These three. Must be making a slight error somewhere. Any help thanks.

Prev Up Next (1 pt) A constant force F -3i 4j 2k m
Prev Up Next (1 pt) A constant force F -3i 4j 2k m
Prev Up Next (1 pt) A constant force F -3i 4j 2k m
Answer #1

The workdone by the force is equal to the dot product of force vector and displacement vector.

The displacement vector is given by

r = (x_2-x_1)i+(y_2-y_1)j+(z_2-z_1)k= (-5-5)i+(3+4)j+(1+2)k \\ = -10i+7j+3k

The work done is

W = F.r = (-3i-4j-2k).(-10i+7j+3k) = 30-28-6 = -4J